Sho w that the following relations hold for a reversible adiabatic 

expansion o f an idea l gas: 

The firebal l o f a uraniu m fissio n bom b consist s o f a spher e o f ga s 

of radiu s 15 m and temperatur e 300000 K shortl y after detonation . 

Assuming that the expansion is adiabatic and that the fireball remain s 

spherical, estimat e th e radiu s o f the bal l whe n th e temperatur e i s 

3000 K. (Tak e y=1.4.) 


1
Expert's answer
2021-04-06T06:30:23-0400

2

these formula will be used


"\bigstar"

TVγ-1 = constant .......(1)


"\bigstar"

T1V1γ-1=T2V2 γ-1 ..........(2)



we can find V1 as we have given radius , and also we know t1 and t2 , further we can find V2


V1 = volume of sphere


V=4/3πr3

(r= 15m)


On solving

We got


V1 = 14137.17m3

and

T1= 300000 K


T2= 3000 K


γ-1 = 1.4 -1

= 0.4

now we solve for V2



V2 γ-1 = "T_1V_1^{γ-1} \over T_2"

On solving we got


V2 = (1413710684.13)m3


And r= "({3V\over 4\pi }){^1\over3}"


r2 = 438.54m